OpenStudy (anonymous):

please help!! 4/5b-4=-9/10b+1

OpenStudy (anonymous):

4/5b+9/10b=5 17/10b=5/*10/17 b=50/17

OpenStudy (anonymous):

oh Thank u !!! u are the first person that has been correct on these questions!